OpenStudy (anonymous):

check please 3√a*5√a i said a ^8/15

OpenStudy (jdoe0001):

\( \large a^{\frac{8}{15}} \ ? \)

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

a to the power of 8/15

OpenStudy (jdoe0001):

are we talking about \( \large 3\sqrt{a}\times 5\sqrt{a} \) ?

OpenStudy (anonymous):

\[\sqrt[3]{a}\times \sqrt[5]{a}\]

OpenStudy (jdoe0001):

ok, then yes, it's \( \large a^{\frac{8}{15}} \)

OpenStudy (anonymous):

alright thank yoooou :)

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

:)